1.9/3+3a/4=13/3
2.5x-2/3=8-3x/5
3.b-2/b+4=b+1/b+10
4.5y-1/2-3y+5=7y+13/6

Answer by AnlytcPhil(1806)

1.

      9/3 + 3a/4 = 13/3

The LCD of the denominators 3, 4, and 3 is 12. So
multiply every term by 12·

12·9/3 + 12·3a/4 = 12·13/3

         36 + 9a = 52

Solve that and get a = 16/9

2.

         5x - 2/3 = 8 - 3x/5

The LCD of the denominators 3 and 5 is 15. So
multiply every term by 15·

   15·5x - 15·2/3 = 15·8 - 15·3x/5

         75x - 10 = 120 - 9x

Solve that and get x = 65/42


3.

         b - 2/b + 4 = b + 1/b + 10

The LCD of the denominators b and b is b. So
multiply every term by 

   b - 2/b + 4 = b + 1/b + 10

         b² - 2 + 4b = b² + 1 + 10b

Solve that and get b = -1/2

4.

         5y - 1/2 - 3y + 5 = 7y + 13/6

I'll let you do that one by yourself. Answer = 7/15

Edwin